A powerful motorcycle can accelerate from rest to 26.8m/s in only 3.8s. How far does it travel in this time? Round two decimal places


1
Expert's answer
2021-02-12T10:43:39-0500
"d=\dfrac{v_0+v}{2}t,""d=\dfrac{0+26.8\ \dfrac{m}{s}}{2}\cdot3.8\ s=50.92\ m."